OpenStudy (anonymous):

Write 4x^3+8x^2-96x in factored form.

OpenStudy (anonymous):

Not sure, might be (x+6)(x-4).

OpenStudy (anonymous):

\[4x ^{3}+8x ^{2}-96x\] To clarify.

OpenStudy (dangerousjesse):

Okay, here's how you'd factorize the equation:

OpenStudy (anonymous):

4 (x-4) x (x+6)

OpenStudy (anonymous):

OK, I knew you had to take out the GCF and factor, but I didn't know what to do with the GCF after that.

OpenStudy (dangerousjesse):

Solve for x over the real numbers: \[4 x (x^2+2 x-24) = 0 \]Divide both sides by a constant to simplify the equation. Divide both sides by 4: \[x (x^2+2 x-24) = 0 \]Solve each term in the product separately. Split into two equations:\[x=0\]or\[ x^2+2 x-24 = 0\]Look at the second equation: Factor the left hand side. The left hand side factors into a product with two terms: \[x = 0 \]or \[(x-4) (x+6) = 0 \]Solve each term in the product separately. Split into two equations:\[x = 0 \]or \[x-4 = 0 \]or\[ x+6 = 0 \]Look at the second equation: Solve for x. Add 4 to both sides:\[x = 0 \]or\[ x = 4 \]or\[ x+6 = 0 \]Look at the third equation: Solve for x. Subtract 6 from both sides:

OpenStudy (anonymous):

x=0, x=4, x=-6?

OpenStudy (dangerousjesse):

Yep

OpenStudy (anonymous):

So that is factored form???

OpenStudy (anonymous):

@DangerousJesse

OpenStudy (dangerousjesse):

Yep

OpenStudy (anonymous):

Oh, ok thanks, that help quite a bit :D

OpenStudy (dangerousjesse):

No problem :)